OpenStudy (minime):

2x+55y=2 3x-2y= -16 please solve by elminaton using full steps

OpenStudy (minime):

hello

OpenStudy (anonymous):

switch both equations so the 2nd equation is on top (it helps) 3x - 2y = -16 2x + 55y = 2 now subtract equation 1 by equation 2 x - 57y = -18 2x 55y = 2 and subtract equation 2 by 2 times equation 1 x - 57y = -18 169y = 38 divide equation 2 by 169 x - 57y = -18 y = 38/169 and then add 57 times equation 2 to equation 1 x = -876/169 y = 38/169

OpenStudy (anonymous):

it would be easier if not for the 55 coefficient. Otherwise you would spend a lot of time double checking for wrong calculations.